calculate standard deviation
X: 2 4. 6. 8. 10
Y: 2. 3. 4. 2. 4
1
Expert's answer
2018-01-24T09:39:08-0500
X: 2 4 6 8 10
Y: 2 3 4 2 4
Standard deviation S = (sum(x - xm)/(n - 1))^0.5, where n = 5, xm is mean.
xmx = (2 + 4 + 6 + 8 + 10)/5 = 6.
Sx = (((2 - 6)^2 + (4 - 6)^2 + (6 - 6)^2 + (8 - 6)^2 + (10 - 6)^2)/4)^0.5 = ((16 + 4 + 0 + 4 + 16)/4)^0.5 = 3.16.
xmy = (2 + 3 + 4 + 2 + 4)/5 = 3.
Sy = (((2 - 3)^2 + (3 - 3)^2 + (4 - 3)^2 + (2 - 3)^2 + (4 - 3)^2)/4)^0.5 = ((1 + 0 + 1 + 1 + 1)/4)^0.5 = 1.